calculate the total no of electron present in 1.6g of of methane

Grade:

1 Answers

Aarti Gupta
9 years ago
GMM (Gram molecular mass) of CH4 =16g
16g of CH4 contain = 6.023 * 1023 electrons
therefore ,
1g CH4 will have = 6.023*1023/16
so,
1.6g of CH4 contain = 6.023*1023/16 * 1.6
=6.023 * 1022 e-